Astronomy Report – February 15th

Crew 205 Astronomy Report 15 Feb 2019

Crew Astronomer: Ghanim Alotaibi

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: No

Objects Viewed: No observational session was requested for tonight. I am waiting for a clear sky for the observations submitted in 13 Feb. We wanted to submit the SY MON photometry measurement to the AAVSO However, some errors were found. The CCD setting was not correct, and there are better options to select comparison stars.

As the new set up was inserted and new comparison stars were selected, the new measured value for the SY MON star was found to be 13.918. However, the magnitude value of the check star was unreasonable and was reported to observatory director.

The crew astronomer used another method to calculate the magnitude value for the target. He measured the instrumental magnitude directly from the software and performed the calculation manually. The magnitude value was found to be 14.118. The value was also reported to observatory director.

Problems Encountered: Nothing to mention

Astronomy Report – February 13th

Crew 205 Astronomy Report 13 Feb 2019

Crew Astronomer: Ghanim Alotaibi

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-14

Objects Viewed: Out of the 18 images requested last night, I found only 12 images were captured. However, all the captured images were found to be bad for photometry. The seeing condition was not good yesterday for astrophotography and it was decided to resubmit the observation session.

Problems Encountered: Bad weather conditions.

Astronomy Report – February 12th

Crew 205 Astronomy Report 12 Feb 2019

Crew Astronomer: Ghanim Alotaibi

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-14

Objects Viewed: ASASSN-V J081823.00-111138.9. I called it MDRS Target 1 in Skynet (RA: 08 18 23.00 , DEC: -11 11 38.9). The observations are planned to be 120 s exposure time with the B filter and 60 s with the V filter. 9 exposures were requested for a period of about 3 hours. A photometry value was calculated for the target SY MON (that was observed in 11 Feb 2019), and the photometry value was found to be 14.147. This value will be submitted to the AAVSO.

Problems Encountered: Nothing to mention

Astronomy Report – February 11th

Crew 205 Astronomy Report 11 Feb 2019

Crew Astronomer: Ghanim Alotaibi

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested:MDRS-14

Objects Viewed: SY MON variable star was requested for observation in Skynet in 10 Feb 2019 night at about 9:15 Local Time. 60 seconds exposure time using V filter was requested for the purpose of photometry.

Problems Encountered: The crew was informed that Skynet is not working properly in the past days due to a software glitch. No Data was obtained the night of 10 Feb 2019. However, this morning we were informed Skynet was nominal working correctly. We expect the observation to have been conducted this evening. We will try to download the data later tonight, and continue conducting photometry tomorrow morning.

Astronomy Report – January 24th

Crew Astronomer: Santiago Vargas Dominguez

Description:

Missing exposure for Horsehead Nebula (Barnard 33) was completed and I processed the complete set of nearly 3 hours in LRGB+Halpha with multiple exposure times (50, 120 and 300 seconds). Please find final cropped image attached.

Objects Viewed: Barnard 33.

Problems Encountered: Dome was close due to the problems with the generator

Astronomy Report – January 23rd

Crew 203 Astronomy Report 23-Jan-2019

Crew Astronomer: Santiago Vargas Dominguez

Description:

The weather improved after a few days under cloudy conditions. This allowed completing most of the active observations from previous runs (Heart Nebula, Barnard 33). Single exposures in LRGB + Halpha (300 s each) were acquired for M82. Images of the moon were acquired for testing with clear skies, but they saturated even with the shortest exposure time (0.03 s). In a previous run, the resulting image of the moon looked fine due to thin clouds blocking some of the light and acting as a neutral filter.

Objects Viewed: Messier 82 (image attached), Soul Nebula, Heart Nebula, Barnard 33.

Problems Encountered: none

Crew 203 Astronomy Report 19Jan2019

[title Astronomy Report – January 19th]

[category astronomy-report]

Hello Mission Support.

Crew 203 Astronomy Report 18-Jan-2019

Crew Astronomer: Santiago Vargas Dominguez

Description: Bad weather in the last couple of nights has given very few chances of pursuing astronomical observations. A set of 100 images of the moon was acquired early this morning (in preparation for the lunar eclipse next sunday) of which only a single frame had moderate quality (away from clouds) to be further preserved (image attached).

Objects Heart Nebula, Barnard 33 are active in the skynet observation list.

Objects Viewed: Moon

Problems Encountered: none

Astronomy Report Jan 18th

Astronomy Report

Crew 203 Astronomy Report 18-Jan-2019

Crew Astronomer: Santiago Vargas Dominguez

Description: Bad weather in the last couple of nights has given very few chances of pursuing astronomical observations. A set of 100 images of the moon was acquired early this morning (in preparation for the lunar eclipse next sunday) of which only a single frame had moderate quality (away from clouds) to be further preserved (image attached).

Objects Heart Nebula, Barnard 33 are active in the skynet observation list.

Objects Viewed: Moon

Problems Encountered: none

Astronomy Report – January 14th

Crew Astronomer: Santiago Vargas Dominguez

Description: Observing run aimed to complement information in Halpha for M31 data (newly processed image attached).

Some data acquired but observation is yet to be completed.

Observation keep active in the skynet observation list.

Objects Viewed: M31, Heart Nebula

Problems Encountered: none